OpenStudy (anonymous):

The National Association for Women in Science asked recent high school grads if they had taken certain science classes. Of those surveyed, 90 said they had taken physics, 138 said they had taken chemistry and 45 said they had taken both. Thirty-two said they had taken neither. How many recent high school grads were surveyed? Show your work

OpenStudy (anonymous):

90 physics 138 chemisty 45 both 32- neither

OpenStudy (anonymous):

i got 146

OpenStudy (anonymous):

how please

OpenStudy (anonymous):

90+138= 228 228/2=114 (the 2 is because 45 took both) and 114+32 ( because 32 took neither) so 146

OpenStudy (anonymous):

thanks